OpenStudy (anonymous):

Find the type and number of solutions for the given equation. 3.2x² - 8.5x + 1.3 = 0

OpenStudy (anonymous):

\[3.2x^2 - 8.5x + 1.3 = 0 \] \[b^2 - 4ac = (-8.5)^2 - 4(3.2)(1.3)\] \[72.25 - 16.64 = 55.61\] @shamil98

OpenStudy (shamil98):

Like i said if the discriminant is greater than zero there are two real solutions, if it is equal to zero then there is one repeating solution , if it negative it has two complex solutions.

OpenStudy (anonymous):

so this one has two real solutions. Awesome! thank you so much!